What is the molality of a 10% by mass NaCl solution when used in qualitative determination of
tannins in plants?
1
Expert's answer
2020-10-08T13:56:48-0400

10% means 100 g of solution contains 10 g of NaCl

M = 58.44

n = 10/58.44 = 0.1711 mol

Density of the solution is 1.09

V = 100 / 1.09 = 91.74 mL = 0.09174 L

Molarity = 0.1711 / 0.09174 = 1.86 M